### v1

@@ 3 EM scheme with exponential family distributions
3.1 Basic properties of EM scheme with exponential family distributions
@@
$p_{\bm{x}}(\bm{\theta})$ $=\beta(\bm{x})\exp\big{(}\bm{\eta}(\bm{\theta})^{\top}\bm{\gamma}(\bm{x})-A(\bm{\theta})\big{)}$ (11)
for some functions $A:\mathbb{R}^{p}\to\mathbb{R}$, $\beta:\Xi\to[0,\infty)$ and vectors of functions $\bm{\eta}:\mathbb{R}^{p}\to\mathbb{R}^{q}$, $\bm{\gamma}:\Xi\to\mathbb{R}^{q}$.
@@
Suppose that any probability distribution $p_{\bm{x}}(\bm{\theta})$ of the exponential family (11) is used for Algorithm 1 and let $\{\bm{\theta}^{(t)}\}_{t\geq 0}$ denote the sequence of solutions generated by the algorithm. Then we have the following statements.
(a) The probability distribution (8) is represented as
$q_{\bm{x}}(\bm{\theta}^{(t-1)})$ $=\tilde{\beta}(\bm{x})\exp\left(\bm{\eta}(\bm{\theta}^{(t-1)})^{\top}\bm{\gamma}(\bm{x})-\tilde{A}(\bm{\theta}^{(t-1)})\right),$ (12)
@@
$\tilde{A}(\bm{\theta}^{(t-1)})$ $:=\ln\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t-1)})}[G(\bm{X})]+A(\bm{\theta}^{(t-1)}),$ (13)
@@
$u(\bm{\theta}|\bm{\theta}^{(t-1)})$ $=-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\big{(}\bm{\eta}(\bm{\theta})-\bm{\eta}(\bm{\theta}^{(t-1)})\big{)}+A(\bm{\theta})-A(\bm{\theta}^{(t-1)})+f(\bm{\theta}^{(t-1)}).$ (14)
@@
(d) Suppose that $\bm{\eta}(\bm{\theta})$ is differentiable and (square and) invertible. Then, we have
$\mathbb{E}_{q_{\bm{x}}(\bm{\theta})}\big{[}\bm{\gamma}(\bm{X})\big{]}$ $=\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]-\big{(}\nabla\bm{\eta}(\bm{\theta})^{\top}\big{)}^{-1}\nabla f(\bm{\theta}),$ (15)
where $\nabla\bm{\eta}(\bm{\theta})$ denotes the Jacobian of $\bm{\eta}(\bm{\theta})$, namely, $\nabla\bm{\eta}(\bm{\theta})=(\nabla\eta_{1}(\bm{\theta}),...,\nabla\eta_{p}(\bm{\theta}))^{\top}$.
Statement (a) shows that when the probability distribution $\big{(}p_{\bm{x}}(\bm{\theta})\big{)}_{\bm{x}}$ is a member of the exponential family, so is the distribution $\big{(}q_{\bm{x}}(\bm{\theta})\big{)}_{\bm{x}}$. Statement (b) states that the lines 3 and 4 of Algorithm 1 can be replaced with
$\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\Theta}\Big{\{}A(\bm{\theta})-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\bm{\eta}(\bm{\theta})\Big{\}}.$ (16)
It also implies that $u(\bm{\theta}|\bm{\theta}^{(t-1)})$ is convex in $\bm{\theta}$ and, thus, (16) is a convex optimization problem if $A(\bm{\theta})-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\bm{\eta}(\bm{\theta})$ is convex in $\bm{\theta}$. Statement (c) yields a sufficient condition for the stationarity of the M-step (16). Statement (d) will be used to derive several formulae, especially, the natural gradient update formula in Theorem 3.3.
@@
3.2 EM scheme with simple exponential family distribution
Building upon the basic properties presented in the previous subsection, we will show that the update formula (16) for a class of exponential family distributions results in natural gradient descent without line search.
Let $\bm{I}_{p_{\bm{x}}}(\bm{\theta})$ denote the Fisher information matrix of a probability distribution $p_{\bm{x}}(\bm{\theta})$, namely,
@@ 3.2 EM scheme with simple exponential family distribution
Suppose that $f$ and $p_{\bm{x}}$ are differentiable and $p_{\bm{x}}(\bm{\theta})$ is a member of the exponential family (11) with $\nabla\bm{\eta}(\bm{\theta})$ being a $p\times p$ invertible matrix, and that there exists an invertible matrix $\bm{\Omega}\in\mathbb{R}^{p\times p}$ such that $\bm{\Omega}\bm{\gamma}(\bm{X})$ is an unbiased estimator of $\bm{\theta}$, namely, it holds that
$\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big{[}\bm{\Omega}\,\bm{\gamma}(\bm{X})\big{]}=\bm{\theta}.$ (17)
@@
$\bm{I}_{p_{\bm{x}}}(\bm{\theta})=\big{(}\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]\big{)}^{-1}=\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1},$ (18)
where $\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\cdot]$ denotes the (co)variance operator under the probability distribution $p_{\bm{x}}(\bm{\theta})$ of $\bm{X}$. In particular, if $f$ is differentiable, the update formula providing the stationarity point of (16) as $\bm{\theta}^{(t)}$ is then reduced to a natural gradient descent of the form:
$\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\bm{I}_{p_{\bm{x}}}(\bm{\theta}^{(t-1)})^{-1}\nabla f(\bm{\theta}^{(t-1)}).$ (19)
@@
Intuitively, the natural gradient (19) modifies the gradient, reflecting the metric of a manifold of the exponential family distribution $p_{\bm{X}}(\bm{\theta})$ via its Fisher information matrix or the covariance matrix of $\bm{\Omega}\bm{\gamma}(\bm{X})$. This comes from the fact that we employ $p_{\bm{X}}(\bm{\theta})$ so as to make the general optimization problem (1) equivalent to minimizing (4). Next section will present three examples to show how an optimization problem can be connected to a natural gradient descent.
4 Three simple examples
@@
4.1 Unconstrained convex quadratic minimization
@@
$\begin{array}[]{ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}&f(\bm{\theta})=\displaystyle\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta},\end{array}$ (21)
where $\bm{Q}\in\mathbb{S}^{p}_{+}$ and $\bm{b}\in\mathbb{R}^{p}$. To apply the EM scheme (Algorithm 1) to solve (21), we employ the $p$-dimensional normal distribution $\bm{{\mathrm{N}}}(\bm{\theta},\bm{\Sigma})$ for random variables $\bm{X}=(X_{1},...,X_{p})$, where $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$. Here, we suppose that the covariance $\bm{\Sigma}$ is known and the mean $\bm{\theta}$ is the only unknown parameter, which corresponds to the decision variables of the problem (21). The normal distribution is a member of the exponential family (11) since its density can be decomposed as
$p_{\bm{x}}(\bm{\theta})$ $=\frac{1}{(2\pi)^{\frac{p}{2}}\sqrt{|\bm{\Sigma}|}}\exp\Big{(}-\frac{1}{2}(\bm{x}-\bm{\theta})^{\top}\bm{\Sigma}^{-1}(\bm{x}-\bm{\theta})\Big{)}$ (22)
$=\underbrace{\exp\big{(}-\frac{1}{2}\bm{x}^{\top}\bm{\Sigma}^{-1}\bm{x}\big{)}}_{\beta(\bm{x})}\exp\big{(}\underbrace{\bm{x}^{\top}}_{\bm{\gamma}(\bm{x})^{\top}}\underbrace{\bm{\Sigma}^{-1}\bm{\theta}}_{\bm{\eta}(\bm{\theta})}\underbrace{-\ln(2\pi)^{\frac{p}{2}}\sqrt{|\bm{\Sigma}|}-\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}}_{-A(\bm{\theta})}\big{)}.$ (23)
@@
Suppose that $\bm{\Sigma}^{-1}\succ\bm{Q}$, where $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$ is the covariance matrix of normal distribution (22), and $\bm{Q}\in\mathbb{S}^{p}_{+}$ is the matrix defining the quadratic objective of (21). If $\bm{X}$ follows the normal distribution (22) and the functions $G$ is defined by
$G(\bm{x})=\sqrt{\frac{|\bm{\Sigma}|}{|\tilde{\bm{Q}}+\bm{\Sigma}^{-1}|}}\exp\left(-\frac{1}{2}\bm{x}^{\top}\tilde{\bm{Q}}\bm{x}-\tilde{\bm{b}}^{\top}\bm{x}\right),$ (24)
where $\tilde{\bm{Q}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\bm{Q})^{-1}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\quad\mbox{and}\quad\tilde{\bm{b}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\bm{Q})^{-1}\bm{b},$ the condition (4) for (21) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (22) and $\Theta=\mathbb{R}^{p}$.
@@
From the correspondence indicated in (23), $\bm{\gamma}(\bm{x})=\bm{x}$ holds for the normal distribution. If we set $\bm{\Omega}=\bm{I}$ for (17), we see $\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big{[}\bm{\Omega}\bm{\gamma}(\bm{X})\big{]}=\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{X}]=\bm{\theta}$, so $\bm{\Omega}\bm{\gamma}(\bm{X})=\bm{\gamma}(\bm{X})$ turns out to be an unbiased estimator of $\bm{X}$. Therefore, the update formula (19) for (16) is simplified as follows.
@@
If the normal distribution (22) is employed for Algorithm 1 to solve (21), the update formula (19) results in
$\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\bm{\Sigma}(\bm{Q}\bm{\theta}^{(t-1)}+\bm{b}).$ (25)
@@
Algorithm 2 EM algorithm for unconstrained convex quadratic minimization (21)
1:Let $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$ satisfy $\bm{\Sigma}^{-1}\succ\bm{Q}$, and $\bm{\theta}^{0}\in\mathbb{R}^{n}$, and set $t\leftarrow 1$.
@@
3:     Update the solution by (25) and $t\leftarrow t+1$.
@@
From the MM optimization perspective, (25) comes from the minimization of the convex quadratic surrogate function
$u(\bm{\theta}|\bar{\bm{\theta}})=\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}-\big{\{}(\bm{\Sigma}^{-1}-\bm{Q})\bar{\bm{\theta}}-\bm{b}\big{\}}^{\top}\big{(}\bm{\theta}-\bar{\bm{\theta}}\big{)}-\frac{1}{2}\bar{\bm{\theta}}^{\top}\bm{\Sigma}^{-1}\bar{\bm{\theta}}+f(\bar{\bm{\theta}}),$ (26)
@@
which can be derived from Proposition 3.1 (b) and whose convexity is stronger than $f(\bm{\theta})$ because of the relation of the Hessian matrices, $\bm{\Sigma}^{-1}\succ\bm{Q}$. The quadratic surrogate function (26), although it was not originally intended to be applied to quadratic optimization in that context, can be considered as a special case of the quadratic lower bound principle [6]. In addition, observing that (26) is decomposed as
$u(\bm{\theta}|\bar{\bm{\theta}})=f(\bm{\theta})+\frac{1}{2}\|\bm{\theta}-\bar{\bm{\theta}}\|^{2}_{\bm{\Sigma}^{-1}-\bm{Q}},$
@@
the MM algorithm can be viewed as a proximal point algorithm with an (elliptical) quadratic proximal term. Recalling that the surrogate functions in our EM scheme are all derived from concavity of the logarithm as in (6), it may be interesting to note that different majorization principles lead to the same surrogate function. While we here assume $\bm{Q}\in\mathbb{S}^{p}_{+}$ so that (21) will be well-defined, the convexity of the quadratic objective is not crucial for the MM scheme. In fact, in Section 5.2 we will consider a possibly nonconvex quadratic objective in developing the EM algorithm for constrained problems with the help of the normal distribution.
@@
To see the role of $\bm{\Sigma}$, assume that $\bm{Q}\in\mathbb{S}^{p}_{++}$ and let $\bm{\Sigma}=\bm{Q}^{-1}$, hypothetically. This case does not satisfy the assumption of Proposition 4.1, but, for such $\bm{\Sigma}$, we could get the update formula as $\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\bm{Q}^{-1}(\bm{Q}\bm{\theta}^{(t-1)}+\bm{b})\equiv-\bm{Q}^{-1}\bm{b}$, which is equivalent to the solution of the problem (21). For a general $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$, the formula (25) can be viewed as a generalization of the steepest descent, i.e., $\bm{\Sigma}=\sigma\bm{I}$ for a constant $\sigma>0$. As a bit more practical example is a diagonal matrix $\bm{\Sigma}=\mathrm{diag}(\sigma_{1},...,\sigma_{p})$ for $\sigma_{j}>0,j\in[p]$. Note that, for example, setting $\sigma_{j}\in(0,1/(q_{jj}+\sum\limits_{h\neq j}|q_{jh}|)),j\in[p],$ makes $\bm{\Sigma}^{-1}-\bm{Q}$ diagonally dominant, thus ensuring positive definite (e.g., [13]). Despite that it permits the step size to vary for each coordinate, the practical advantage of this modification over the steepest descent may not be as significant as one might expect. However, the relationship between the convex quadratic objective and the normal distribution described above can be used for the example of Section 4.1.1 below.
@@
Convergence rate can be shown by applying existing results on EM algorithms [19, 18].

### v2

@@ 3. EM scheme with exponential family distributions
3.1. Basic properties of EM scheme with exponential family distributions
@@
$p_{\bm{x}}(\bm{\theta})$ $=\beta(\bm{x})\exp\big({\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{\theta})}-A(\bm{\theta})\big)$ (3.1)
for some functions $A:\mathbb{R}^{p}\to\mathbb{R}$, $\beta:\Xi\to[0,\infty)$ and vectors of functions $\bm{\eta}:\mathbb{R}^{p}\to\mathbb{R}^{q}$, $\bm{\gamma}:\Xi\to\mathbb{R}^{q}$. The exponential family encompasses a wide array of representative distributions (see, for instance, [8]). Subsequent sections will demonstrate the significance of certain exponential family distributions, including the normal distribution, binomial distribution, multinomial distribution, and Poisson distribution, in the context of the relationship between EM schemes and polynomial objective optimization problems. First, the basic properties of the exponential family when applied to Algorithm 1 are summarized in the following proposition.
@@
Suppose that any probability distribution $p_{\bm{x}}(\bm{\theta})$ of the exponential family (3.1) is used for Algorithm 1 and let $\{\bm{\theta}^{(t)}\}_{t\geq 0}$ denote the sequence of solutions generated by the algorithm. Then we have the following statements.
(a) The probability distribution (2.8) is represented as
$q_{\bm{x}}(\bm{\theta}^{(t-1)})$ $=\tilde{\beta}(\bm{x})\exp\left(\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{\theta}^{(t-1)})-\tilde{A}(\bm{\theta}^{(t-1)})\right),$ (3.2)
@@
$\tilde{A}(\bm{\theta}^{(t-1)})$ $:=\ln\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t-1)})}[G(\bm{X})]+A(\bm{\theta}^{(t-1)}),$ (3.3)
@@
$u(\bm{\theta}|\bm{\theta}^{(t-1)})$ $=-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\big(\bm{\eta}(\bm{\theta})-\bm{\eta}(\bm{\theta}^{(t-1)})\big)+A(\bm{\theta})-A(\bm{\theta}^{(t-1)})+f(\bm{\theta}^{(t-1)}).$ (3.4)
@@
(d) Suppose that $\bm{\eta}(\bm{\theta})$ is square (i.e., $q=p$) and invertible, and each component $\eta_{i}(\bm{\theta})$ is differentiable, $i\in[p](=[q])$. Then, we have
$\mathbb{E}_{q_{\bm{x}}(\bm{\theta})}\big[\bm{\gamma}(\bm{X})\big]$ $=\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{\gamma}(\bm{X})]-\big(\nabla\bm{\eta}(\bm{\theta})^{\top}\big)^{-1}\nabla f(\bm{\theta}),$ (3.5)
where $\nabla\bm{\eta}(\bm{\theta})$ denotes the Jacobian of $\bm{\eta}(\bm{\theta})$, i.e., $\nabla\bm{\eta}(\bm{\theta}):=(\nabla\eta_{1}(\bm{\theta}),...,\nabla\eta_{p}(\bm{\theta}))^{\top}$.
Statement (a) shows that when the probability distribution $\big(p_{\bm{x}}(\bm{\theta})\big)_{\bm{x}}$ is a member of the exponential family, so is the distribution $\big(q_{\bm{x}}(\bm{\theta})\big)_{\bm{x}}$. Statement (b) states that lines 3 and 4 of Algorithm 1 can be replaced with
$\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\Theta}\Big\{A(\bm{\theta})-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\bm{\eta}(\bm{\theta})\Big\}.$ (3.6)
It also implies that $u(\bm{\theta}|\bm{\theta}^{(t-1)})$ is convex in $\bm{\theta}$ and, thus, (3.6) is a convex optimization problem if $A(\bm{\theta})-\mathbb{E}_{q_{\bm{x}}(\bm{\theta}^{(t-1)})}[\bm{\gamma}(\bm{X})]^{\top}\bm{\eta}(\bm{\theta})$ is convex in $\bm{\theta}$. Statement (c) yields a sufficient condition for the stationarity of the M-step (3.6). Statement (d) will be used to derive several formulae, especially, the natural gradient update formula in Theorem 3.3.
@@
3.2. EM scheme with simple exponential family distribution
Building upon the basic properties presented in the previous subsection, we will show that the update formula (3.6) for a class of exponential family distributions results in natural gradient descent without line search.
Let $\bm{I}_{p_{\bm{x}}}(\bm{\theta})$ denote the Fisher information matrix of a probability distribution $p_{\bm{x}}(\bm{\theta})$, i.e.,
@@ 3.2. EM scheme with simple exponential family distribution
Suppose that $p_{\bm{x}}(\bm{\theta})$ is differentiable and a member of the exponential family (3.1) with $\nabla\bm{\eta}(\bm{\theta})$ being a $p\times p$ invertible matrix, and that there exists an invertible matrix $\bm{\Omega}\in\mathbb{R}^{p\times p}$ such that $\bm{\Omega}\bm{\gamma}(\bm{X})$ is an unbiased estimator of $\bm{\theta}$, i.e.,
$\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big[\bm{\Omega}\,\bm{\gamma}(\bm{X})\big]=\bm{\theta}.$ (3.7)
@@
$\bm{I}_{p_{\bm{x}}}(\bm{\theta})=\big(\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\bm{\Omega}\bm{\gamma}(\bm{X})]\big)^{-1}=\nabla\bm{\eta}(\bm{\theta})^{\top}\bm{\Omega}^{-1},$ (3.8)
where $\mathbb{V}_{p_{\bm{x}}(\bm{\theta})}[\cdot]$ denotes the (co)variance operator under the probability distribution $p_{\bm{x}}(\bm{\theta})$ of $\bm{X}$. In particular, if $f$ is differentiable, the update formula providing the stationarity point of (3.6) is then reduced to a natural gradient descent of the form:
$\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\bm{I}_{p_{\bm{x}}}(\bm{\theta}^{(t-1)})^{-1}\nabla f(\bm{\theta}^{(t-1)}).$ (3.9)
@@
Intuitively, the natural gradient (3.9) modifies the gradient, reflecting the metric of a manifold of the exponential family distribution $p_{\bm{x}}(\bm{\theta})$ via its Fisher information matrix or the covariance matrix of $\bm{\Omega}\bm{\gamma}(\bm{X})$. This comes from the fact that we employ $p_{\bm{x}}(\bm{\theta})$ so as to make the general optimization problem (2.1) equivalent to minimizing (2.4). Next section will present three examples to show how an optimization problem can be connected to a natural gradient descent.
Remark 3.4.
Theorem 3.3 assumes that $\bm{\Omega}$ does not depend on $\bm{\theta}$. If, on the other hand, $\bm{\Omega}$ depends on $\bm{\theta}$, i.e., $\bm{\Omega}=\bm{\Omega}(\bm{\theta})$, we can see that (3.9) becomes somewhat more complicated, such as
$\bm{\theta}^{(t)}\leftarrow\bm{\Omega}(\bm{\theta}^{(t)})\bm{\Omega}(\bm{\theta}^{(t-1)})^{-1}\big(\bm{\theta}^{(t-1)}-\bm{I}_{p_{\bm{x}}}(\bm{\theta}^{(t-1)})^{-1}\nabla f(\bm{\theta}^{(t-1)})\big).$
All the examples demonstrated below are constructed based on the constant matrix $\bm{\Omega}$, and as a result we can enjoy the simple formula (3.9).
4. Three simple examples
@@
4.1. Unconstrained convex quadratic minimization
@@
$\begin{array}[]{ll}\underset{\bm{\theta}\in\mathbb{R}^{p}}{\mbox{minimize}}\quad&f(\bm{\theta})=\displaystyle\frac{1}{2}\bm{\theta}^{\top}\bm{Q}\bm{\theta}+\bm{b}^{\top}\bm{\theta},\end{array}$ (4.2)
where $\bm{Q}\in\mathbb{S}^{p}_{+}$ and $\bm{b}\in\mathbb{R}^{p}$. To apply the EM scheme (Algorithm 1) to solve (4.2), we employ the $p$-dimensional normal distribution $\bm{{\mathrm{N}}}(\bm{\theta},\bm{\Sigma})$ for random variables $\bm{X}=(X_{1},...,X_{p})$, where $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$. Here, we suppose that the covariance $\bm{\Sigma}$ is known and the mean $\bm{\theta}$ is the only unknown parameter, which corresponds to the decision variables of the problem (4.2). The normal distribution is a member of the exponential family (3.1) since its density can be decomposed as
$p_{\bm{x}}(\bm{\theta})$ $=\frac{1}{(2\pi)^{\frac{p}{2}}\sqrt{|\bm{\Sigma}|}}\exp\Big(-\frac{1}{2}(\bm{x}-\bm{\theta})^{\top}\bm{\Sigma}^{-1}(\bm{x}-\bm{\theta})\Big)$ (4.3)
$=\underbrace{\exp\big(-\frac{1}{2}\bm{x}^{\top}\bm{\Sigma}^{-1}\bm{x}\big)}_{=\beta(\bm{x})}\exp\big(\underbrace{\bm{x}^{\top}\bm{\Sigma}^{-1}\bm{\theta}}_{=\bm{\gamma}(\bm{x})^{\top}\bm{\eta}(\bm{\theta})}\underbrace{-\ln(2\pi)^{\frac{p}{2}}\sqrt{|\bm{\Sigma}|}-\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}}_{=-A(\bm{\theta})}\big).$ (4.4)
@@
Suppose that $\bm{\Sigma}^{-1}\succ\bm{Q}$, where $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$ is the covariance matrix of normal distribution (4.3), and $\bm{Q}\in\mathbb{S}^{p}_{+}$ is the matrix defining the quadratic objective of (4.2). If $\bm{X}$ follows the normal distribution (4.3) and the functions $G$ is defined by
$G(\bm{x})=\sqrt{\frac{|\bm{\Sigma}|}{|\tilde{\bm{Q}}+\bm{\Sigma}^{-1}|}}\exp\left(-\frac{1}{2}\bm{x}^{\top}\tilde{\bm{Q}}\bm{x}-\tilde{\bm{b}}^{\top}\bm{x}\right),$ (4.5)
where $\tilde{\bm{Q}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\bm{Q})^{-1}\bm{\Sigma}^{-1}-\bm{\Sigma}^{-1}\quad\mbox{and}\quad\tilde{\bm{b}}:=\bm{\Sigma}^{-1}(\bm{\Sigma}^{-1}-\bm{Q})^{-1}\bm{b},$ the condition (2.4) for (4.2) is fulfilled with $p_{\bm{x}}(\bm{\theta})$ defined by (4.3) and $\Theta=\mathbb{R}^{p}$.
@@
From the correspondence indicated in (4.4), $\bm{\gamma}(\bm{x})=\bm{x}$ holds for the normal distribution. If we set $\bm{\Omega}=\bm{I}$ for (3.7), we see $\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big[\bm{\Omega}\bm{\gamma}(\bm{X})\big]=\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}[\bm{X}]=\bm{\theta}$, so $\bm{\Omega}\bm{\gamma}(\bm{X})=\bm{\gamma}(\bm{X})$ turns out to be an unbiased estimator of $\bm{X}$. Therefore, the update formula (3.9) for (3.6) is simplified as follows.
@@
If the normal distribution (4.3) is employed for Algorithm 1 to solve (4.2), the update formula (3.9) results in
$\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\bm{\Sigma}(\bm{Q}\bm{\theta}^{(t-1)}+\bm{b}).$ (4.6)
@@
Algorithm 2 EM algorithm for unconstrained convex quadratic minimization (4.2)
1:Let $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$ satisfy $\bm{\Sigma}^{-1}\succ\bm{Q}$, and $\bm{\theta}^{(0)}\in\mathbb{R}^{n}$, and set $t\leftarrow 1$.
@@
3:  Update the solution by (4.6) and $t\leftarrow t+1$.
@@
From the MM optimization perspective shown in Section 2, the natural gradient formula (4.6) comes from the minimization of the convex quadratic surrogate function
$u(\bm{\theta}|\bm{\theta}^{(t-1)})$ $=\frac{1}{2}\bm{\theta}^{\top}\bm{\Sigma}^{-1}\bm{\theta}-\big\{(\bm{\Sigma}^{-1}-\bm{Q})\bm{\theta}^{(t-1)}-\bm{b}\big\}^{\top}\big(\bm{\theta}-\bm{\theta}^{(t-1)}\big)$
$\quad\qquad-\frac{1}{2}(\bm{\theta}^{(t-1)})^{\top}\bm{\Sigma}^{-1}\bm{\theta}^{(t-1)}+f(\bm{\theta}^{(t-1)}),$ (4.7)
@@
which can be derived from Proposition 3.1 (b) and whose convexity is stronger than $f(\bm{\theta})$ because of the relation of the Hessian matrices, $\nabla^{2}u(\bm{\theta}|\bar{\bm{\theta}})=\bm{\Sigma}^{-1}\succ\bm{Q}=\nabla^{2}f(\bm{\theta})$. The quadratic surrogate function (4.7), although it was not originally intended to be applied to quadratic optimization in that context, can be considered as a special case of the quadratic lower bound principle [9]. In addition, observing that (4.7) is decomposed as
$u(\bm{\theta}|\bm{\theta}^{(t-1)})=f(\bm{\theta})+\frac{1}{2}\|\bm{\theta}-\bm{\theta}^{(t-1)}\|^{2}_{\bm{\Sigma}^{-1}-\bm{Q}},$
@@
the MM algorithm can be viewed as a proximal point algorithm with an (elliptical) quadratic proximal term. Recalling that the surrogate functions in our EM scheme are all derived from concavity of the logarithm as in (2.6), it may be interesting to note that different majorization principles lead to the same surrogate function.
@@
To see the role of the covariance matrix $\bm{\Sigma}$ in the formula (4.6), assume that $\bm{Q}\in\mathbb{S}^{p}_{++}$ and let $\bm{\Sigma}=\bm{Q}^{-1}$, hypothetically. This case does not satisfy the assumption of Proposition 4.1, but, for such $\bm{\Sigma}$, we could get the update formula as $\bm{\theta}^{(t)}\leftarrow\bm{\theta}^{(t-1)}-\bm{Q}^{-1}(\bm{Q}\bm{\theta}^{(t-1)}+\bm{b})=-\bm{Q}^{-1}\bm{b}$, which is equivalent to the solution of the problem (4.2). For a general $\bm{\Sigma}\in\mathbb{S}^{p}_{++}$, the formula (4.6) can be viewed as a generalization of the steepest descent, i.e., $\bm{\Sigma}=\sigma\bm{I}$ for a constant $\sigma>0$. As a bit more practical example is a diagonal matrix $\bm{\Sigma}=\mathrm{diag}(\sigma_{1},...,\sigma_{p})$ for $\sigma_{j}>0,j\in[p]$. Note that, for example, setting $\sigma_{j}\in(0,1/(q_{jj}+\sum_{h\neq j}|q_{jh}|)),j\in[p],$ makes $\bm{\Sigma}^{-1}-\bm{Q}$ diagonally dominant, thus ensuring positive definite (e.g., [16]). Despite that it permits the step size to vary for each coordinate, the practical advantage of this modification over the steepest descent may not be as significant as one might expect. However, the relationship between the convex quadratic objective and the normal distribution described above can be used for the example of Section 4.1.1 below. Note also that while we here assume $\bm{Q}\in\mathbb{S}^{p}_{+}$ so that (4.2) will be well-defined, the convexity of the quadratic objective is not crucial for the MM scheme. In fact, in Section 5.2 we will consider a possibly nonconvex quadratic objective in developing the EM algorithm for constrained problems with the help of the normal distribution.
@@
Convergence rate of Algorithm 2 can be shown by applying existing results on EM algorithms [22, 21].